# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for radiance field rendering, but it typically requires millions of redundant Gaussian primitives, overwhelming memory and rendering budgets. Existing compaction approaches address this by pruning Gaussians based on heuristic importance scores, without global fidelity guarantee. To bridge this gap, we propose a novel optimal transport perspective that casts 3DGS compaction as global Gaussian mixture reduction. Specifically, we first minimize the composite transport divergence over a KD-tree partition to produce a compact geometric representation, and then decouple appearance from geometry by fine-tuning color and opacity attributes with far fewer Gaussian primitives. Experiments on benchmark datasets show that our method (i) yields negligible loss in rendering quality (PSNR, SSIM, LPIPS) compared to vanilla 3DGS with only 10% Gaussians; and (ii) consistently outperforms state-of-the-art 3DGS compaction techniques. Notably, our method is applicable to any stage of vanilla or accelerated 3DGS pipelines, providing an efficient and agnostic pathway to lightweight neural rendering.
@@ 1 Introduction
Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media [1, 2]. The current state-of-the-art, 3D Gaussian Splatting (3DGS) [3], iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes. During rendering, these 3D Gaussians are projected to 2D screens and $\alpha$-blended to achieve real-time photorealistic synthesis.
However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes [4, 5, 6]. This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices [7].
@@
A common strategy to improve efficiency is compaction [8, 9, 10, 7]—reducing the number of Gaussians while preserving rendering fidelity. With fewer primitives, both storage requirements and rendering efficiency (particularly per-frame rendering time) are naturally improved. This approach is viable because 3DGS densification inherently generates redundant primitives [4, 5, 6]. Existing methods achieve compaction via pruning or random subset selection [11, 12, 5, 13, 14, 15, 6, 4, 16]. These strategies, which are based on naïvely discarding information, are simple to implement, but may be ineffective. In particular, they tend to lose critical structural details (as shown in Figure 1) or distort the underlying geometry (as shown in Figure 2). Such losses can degrade rendering quality, particularly in regions with fine-scale features or complex material properties. These limitations motivate our key question:
@@
For GMR, we minimize the composite transportation divergence [17], which is rooted in optimal transport theory [18, 19] and allows an effective algorithm. We tailor this GMR algorithm (detailed in Section 3.2.1) so that it scales well in scenes like 3DGS with an extensive amount of Gaussians. Crucially, our GMR optimizer does not merely select a subset of existing Gaussians; instead, it creates new primitives that can dynamically adjust their positions and covariances to better approximate the underlying geometry (as shown in Figure 2). Our method is fundamentally algorithm-agnostic: it functions as a plug-and-play module that can enhance both the standard 3DGS pipeline and any of its variants (Section 4), applicable at any stage of training to boost computational efficiency.
@@
  • We open a new pathway to view 3DGS representations as a Gaussian mixture and perform compaction from the perspective of Gaussian mixture reduction via optimal transport.
  • We decouple geometry and appearance optimization in a scalable blockwise optimization scheme.
  • Empirical results demonstrate that our method preserves rendering quality using only 10% Gaussians.
  • We provide a plug-and-play module compatible with any 3DGS variant at any training stage.
@@ 2 Related Works
Compaction Techniques in 3DGS 3DGS compaction seeks to minimize the Gaussian count while preserving image quality. Existing work falls into two operations: densification (where to add) and pruning (what to drop). Most current approaches rely on per-Gaussian heuristic scores.
For densification, Taming-3DGS [10] ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS [20] considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro [9] guides growth using depth and normal maps. In the pruning phase, LightGaussian [11], Mini-Splatting [12], RadSplat [5], and AtomGS [13] compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians. Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) [14] or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) [15]. The score can also be trained via a learnable mask, as in LP-3DGS [6], Compact3DGS [4], and HAC [16]. Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes [21] or visible only in real but not virtual views [22]. For a comprehensive survey of 3DGS compression and compaction, we refer readers to [8, 23].
@@
3DGS from Probabilistic Distribution Point of View Kheradmand et al. [24] formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information. However, their update remains pairwise and lacks convergence guarantees under a principled divergence. Moreover, this scheme is coupled with the original 3DGS pipeline, limiting its generalizability to other variants.
@@ 3 Method: Gaussian Herding across Pens
Figure 3: An illustration of the proposed GHAP-3DGS approach. The process begins with full-resolution 3DGS training to obtain initial geometric and appearance features. These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks—analogous to sheep pens. We then perform blockwise Gaussian Model Reduction (GMR) to generate a small number of representative Gaussians within each block. This step is analogous to the popular kernel herding method  [25]. Finally, a lightweight appearance refinement step further optimizes the appearance feature of the reduced set. This multi-stage pipeline progressively guides the Gaussians in each block—analogous to herding across pens—toward a compact and high-fidelity representation.
We present a principled approach for compacting 3DGS representations through Gaussian Mixture Reduction (GMR). Our key insight is that The 3DGS densification strategy, while effective for view synthesis, produces redundant Gaussian components that can be consolidated without sacrificing visual quality. We achieve this by interpreting the 3D scene surface as an unnormalized probability density function (PDF), where the trained 3DGS representation serves as a Gaussian mixture approximation of this PDF. Thus, our problem reduces to finding a Gaussian mixture with fewer components to approximate the original, redundant one. Our procedure separately optimizing geometric (position $\mu$, covariance $\Sigma$) and appearance (opacity $\alpha$, color $c$) attributes to ensure rendering fidelity. The key idea of our method is illustrated in Figure 3.
@@ 3.1 Probabilistic Scene Representation
[…]
@@ 3.2 Compaction via Optimal Transport
Motivated by the observation that many 3DGS algorithms [4, 5, 6] produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity. The process consists of two key phases:
@@
    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) [26], approximating the original Gaussian mixture with redundant components with fewer components. This yields a compacted geometric representation: $\bar{\phi}_{m}(x)=\sum_{j=1}^{m}\bar{\alpha}_{j}\phi(x;\bar{\mu}_{j},\bar{\Sigma}_{j})$, where $m\ll n$. This step modifies only the Gaussian positions ($\bar{\mu}_{j}$) and covariances ($\bar{\Sigma}_{j}$), leaving appearance attributes unchanged.
@@ 3.2.1 Geometric Compaction via GMR
Following Zhang et al. [17], we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
@@
The CTD generalizes optimal transport [18] to mixtures, treating each component as a discrete distribution in Gaussian space. The cost function measures the cost of moving one unit of Gaussian from one location to another, and $\pi_{ij}$ measures the corresponding amount of mass that is being moved. The total cost is proportional to the cost and the mass, and the divergence is the smallest transportation cost to move the original mixture to the target mixture. The reduced mixture becomes
@@
As shown in Zhang et al. [17], (1) can be solved using the effective iterative algorithm in Algorithm 1. The algorithm reduces to a $k$-means variant in Gaussian space: 1) The assignment step follows the same principle as traditional $k$-means, but replaces the $L^{2}$ distance between vectors with a cost function $c(\cdot,\cdot)$ between Gaussian distributions. 2) The update step generalizes the cluster center computation: In traditional $k$-means, centers are updated as arithmetic averages (barycenters w.r.t. $L^{2}$ distance) of vectors in each cluster. In this algorithm, centers become barycenters of Gaussians in each cluster, minimized w.r.t. cost function $c(\cdot,\cdot)$. Thus, standard $k$-means emerges as a special case when using $L^{2}$ distance on vectorized Gaussian parameters.
While the standard GMR algorithm provides optimal theoretical guarantees, its direct application to 3DGS compaction proves computationally prohibitive. Although the algorithm must converge in finite steps [17] [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration. In typical 3D scenes, the number of Gaussians scales as $n=\Omega(10^{5})$ [^3] The $a_{n}=\Omega(b_{n})$ ($O(b_{n})$) if there exist $C\geq 0$ such that $a_{n}\geq Cb_{n}$ ($a_{n}\leq Cb_{n}$)., and even after 95% reduction, each iteration would still require at least $10^{8}$ operations and memory storage. For the update step, the computational cost for the cluster center depends on the pre-specified cost function $c(\cdot,\cdot)$. The KL divergence considered in Zhang et al. [17] suffers from its key limitations: it requires computing inverse covariance matrices, which introduces (1) numerical instability due to small eigenvalues of covariance matrices, and (2) significant overhead despite operating in ${\mathbb{R}}^{3}$ (with $O(\rho s^{2}\log n)$ inversions needed). To overcome this challenge, we introduce two key optimizations designed for 3DGS:
  • Blockwise GMR via KD-Tree: To improve computational and memory efficiency during training, we partition the scene into spatially localized blocks and perform GMR independently within each block. As demonstrated in Remark 1, this blockwise approach yields significant computational savings. While both KD-trees [27] and Octrees [28] are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages. First, it produces more balanced partitions across regions. Second, it avoids unnecessary subdivisions in sparse regions that would waste computational resources.
@@
  • Efficient Cost Function: We propose a novel cost function that is both computationally efficient and maintains approximation quality, overcoming limitations of KL divergence used in [17]. We propose to use the following divergence:
@@
    which has three key advantages: First, it preserves distributional similarity as Gaussians are uniquely determined by their mean and covariance. Second, the assignment step requires only efficient vector/matrix norm computations. Third, the update step simplifies to weighted averages:
    $\bar{\mu}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}.$
3.2.2 Appearance Optimization
@@
4:1. Build KD-tree from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$
5:2. For each leaf block ${\mathcal{B}}_{k}$:
@@
9:$\bar{c}_{j}\leftarrow c_{i^{*}}$ and $\bar{\alpha}_{j}\leftarrow\alpha_{i^{*}}$ where $i^{*}=\operatorname*{arg\,min}_{i\in[n]}\|\mu_{i}-\bar{\mu}_{j}\|_{2}$
@@
Following geometric compaction, we initialize the appearance attributes (opacity and colors) of the compacted Gaussian primitives. For each primitive in the reduced mixture, we assign the appearance parameters from its closest counterpart in the original Gaussian mixture. Using these initial values, we then optimize the appearance attributes through backpropagation following the standard 3DGS training pipeline. We optimize the opacity instead of directly using the opacity from the GMR algorithm since these output weights not necessarily satisfy the constraint on the opacity that it must be within 00 and $1$. Finetuning the the opacity can achieve better visualization performance.

### v2

# Gaussian Herding across Pens:: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for radiance field rendering, but it typically requires millions of redundant Gaussian primitives, overwhelming memory and rendering budgets. Existing compaction approaches address this by pruning Gaussians based on heuristic importance scores, without global fidelity guarantee. To bridge this gap, we propose a novel optimal transport perspective that casts 3DGS compaction as global Gaussian mixture reduction. Specifically, we first minimize the composite transport divergence over a KD-tree partition to produce a compact geometric representation, and then decouple appearance from geometry by fine-tuning color and opacity attributes with far fewer Gaussian primitives. Experiments on benchmark datasets show that our method (i) yields negligible loss in rendering quality (PSNR, SSIM, LPIPS) compared to vanilla 3DGS with only 10% Gaussians; and (ii) consistently outperforms state-of-the-art 3DGS compaction techniques. Notably, our method is applicable to any stage of vanilla or accelerated 3DGS pipelines, providing an efficient and agnostic pathway to lightweight neural rendering. The code is publicly available at https://github.com/DrunkenPoet/GHAP
@@ 1 Introduction
Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media (fei20243d, ; chen2024survey, ). The current state-of-the-art, 3D Gaussian Splatting (3DGS) (kerbl20233d, ), iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes. During rendering, these 3D Gaussians are projected to 2D screens and $\alpha$-blended to achieve real-time photorealistic synthesis.
However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ). This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices (ren2024octree, ).
@@
A common strategy to improve efficiency is compaction (bagdasarian20243dgs, ; cheng2024gaussianpro, ; mallick2024taming, ; ren2024octree, )—reducing the number of Gaussians while preserving rendering fidelity. Fewer primitives result in reduced storage needs and faster rendering, improving per-frame performance. This approach is viable because 3DGS densification inherently generates redundant primitives (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ). Existing methods achieve compaction via pruning or random subset selection (fan2024lightgaussian, ; fang2024mini, ; niemeyer2024radsplat, ; liu2024atomgs, ; ali2024trimming, ; hanson2024pup, ; zhang2024lp, ; lee2024compact, ; chen2024hac, ). These strategies naïvely discard Gaussian primitives; while simple to implement, they are often ineffective. In particular, they tend to lose critical structural details (as shown in Fig. 1) or distort the underlying geometry (as shown in Fig. 2). Such losses can degrade rendering quality, particularly in regions with fine-scale features or complex material properties. These limitations motivate our key question:
@@
For GMR, we minimize the composite transportation divergence (zhang2023gaussian, ), which is rooted in optimal transport theory (peyre2019computational, ; villani2009optimal, ) and allows an effective algorithm. We tailor this GMR algorithm (detailed in Section 3.2.1) so that it scales well in scenes like 3DGS with an extensive amount of Gaussians. Crucially, our GMR optimizer does not merely select a subset of existing Gaussians; instead, it creates new primitives that can dynamically adjust their positions and covariances to better approximate the underlying geometry (as shown in Figure 2). Our method is fundamentally algorithm-agnostic: it functions as a plug-and-play module that can enhance both the standard 3DGS pipeline and any of its variants (Section 4), applicable at any stage of training to boost computational efficiency.
@@
  • We open a new pathway to view 3DGS representations as a Gaussian mixture and perform compaction from the perspective of Gaussian mixture reduction via optimal transport. This contrasts with prior compaction methods that ignore geometric structure and often produce distortions, whereas our approach preserves geometric fidelity, offering a new and impactful direction for 3DGS.
  • We are the first to adapt GMR to 3DGS. We introduce a novel cost function that yields closed-form, low-cost updates. We also develop a block-wise GMR algorithm guided by a KD-tree, enabling efficient large‑scale scene compaction. These strategies are non-trivial and bridges theory with practical scalability.
  • Our method is post-hoc and compatible with any existing 3DGS pipeline, making it highly practical and broadly applicable. With minimal overhead, our approach achieves SOTA compaction performance, both in quality and efficiency.
  • Empirical results demonstrate that our method preserves rendering quality at 10% retention ratio.
@@ 2 Related Works
Compaction Techniques in 3DGS. 3DGS compaction seeks to minimize the Gaussian count while preserving image quality. Existing work falls into two operations: densification (where to add) and pruning (what to drop). Most current approaches rely on per-Gaussian heuristic scores.
For densification, Taming‑3DGS (mallick2024taming, ) ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS (kim2024color, ) considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro (cheng2024gaussianpro, ) guides growth using depth and normal maps. In the pruning phase, LightGaussian (fan2024lightgaussian, ), Mini-Splatting (fang2024mini, ), RadSplat (niemeyer2024radsplat, ), and AtomGS (liu2024atomgs, ) compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians. Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) (ali2024trimming, ) or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) (hanson2024pup, ). The score can also be trained via a learnable mask, as in LP-3DGS (zhang2024lp, ), Compact3DGS (lee2024compact, ), and HAC (chen2024hac, ). Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes (matsuki2024gaussian, ) or visible only in real but not virtual views (ji2024neds, ). For a comprehensive survey of 3DGS compression and compaction, we refer readers to (bagdasarian20243dgs, ; bao20253d, ).
@@
3DGS from Probabilistic Distribution Point of View. Kheradmand et al. (kheradmand20243d, ) formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information. However, their update remains pairwise and lacks convergence guarantees under a principled divergence. Moreover, this scheme is coupled with the original 3DGS pipeline, limiting its generalizability to other variants.
@@ 3 Method: Gaussian Herding across Pens
Figure 3: An illustration of the proposed GHAP approach. The process begins with full-resolution 3DGS training to obtain initial geometric and appearance features. These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks–analogous to sheep pens. We then perform blockwise Gaussian Mixture Reduction (GMR) to approximate the geometric shape within each block using a much smaller number of Gaussians. This step is analogous to the popular kernel herding method  (DBLP:journals/corr/abs-1203-3472, ). Finally, a lightweight appearance refinement step further optimizes the appearance feature of the reduced set. This multi-stage pipeline progressively guides the Gaussians in each block–analogous to herding across pens–toward a compact and high-fidelity representation.
@@ 3.1 Probabilistic Scene Representation
where $\alpha_{i}$ are the mixture weights satisfying $\alpha_{i}>0$ [^1] In statistics, Gaussian mixtures are defined with the constraint $\sum_{i=1}^{n}\alpha_{i}=1$ to ensure the area under the density function is $1$. However, we relax this constraint and consider unnormalized Gaussian mixtures, as in 3DGS, where the integral under the geometric surface need not equal $1$.. In the context of 3DGS, let $x\in{\mathbb{R}}^{3}$ be spatial coordinates and $f(x)$ represent the implicit surface function (i.e., geometric shape and opacity that excludes color). The training process in 3DGS learns opacity parameter $\alpha_{i}$, location parameter $\mu_{i}$, and shape parameter $\Sigma_{i}$ such that
@@ 3.2 Compaction via Optimal Transport
Motivated by the observation that many 3DGS algorithms (lee2024compact, ; niemeyer2024radsplat, ; zhang2024lp, ) produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity. The process consists of two key phases:
@@
    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) (crouse2011look, ), approximating the original Gaussian mixture with redundant components by one with fewer components. This yields a compacted geometric representation: $\bar{\phi}_{m}(x)=\sum_{j=1}^{m}\bar{\alpha}_{j}\phi(x;\bar{\mu}_{j},\bar{\Sigma}_{j})$, where $m\ll n$. This step modifies only the Gaussian positions ($\bar{\mu}_{j}$) and covariances ($\bar{\Sigma}_{j}$), leaving appearance attributes unchanged.
@@ 3.2.1 Geometric Compaction via GMR
Following Zhang et al. (zhang2023gaussian, ), we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
@@
The CTD generalizes optimal transport (peyre2019computational, ) to mixtures, treating each component as a discrete distribution in the space of Gaussian distributions. The cost function measures the cost of moving one unit of Gaussian from one location to another, and $\pi_{ij}$ measures the corresponding amount of mass that is being moved. The total cost is proportional to the cost and the mass, and the divergence is the smallest transportation cost to move the original mixture to the target mixture. The reduced mixture becomes
@@
As shown in Zhang et al. (zhang2023gaussian, ), (1) can be solved using the effective iterative algorithm in Algorithm 1.
@@
The algorithm reduces to a $k$-means variant in Gaussian space: 1) The assignment step follows the same principle as traditional $k$-means, but replaces the $L^{2}$ distance between vectors with a cost function $c(\cdot,\cdot)$ between Gaussian distributions. 2) The update step generalizes the cluster center computation: In traditional $k$-means, centers are updated as arithmetic averages (barycenters w.r.t. $L^{2}$ distance) of vectors in each cluster. In this algorithm, centers become barycenters of Gaussians in each cluster, minimized w.r.t. cost function $c(\cdot,\cdot)$. Thus, standard $k$-means emerges as a special case when using $L^{2}$ distance on vectorized Gaussian parameters.
While the standard GMR algorithm provides optimal theoretical guarantees, its direct application to 3DGS compaction proves computationally prohibitive. Although the algorithm must converge in finite steps (zhang2023gaussian, ) [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration. In typical 3D scenes, the number of Gaussians scales as $n=\Omega(10^{5})$ [^3] The $a_{n}=\Omega(b_{n})$ ($O(b_{n})$) if there exist $C\geq 0$ such that $a_{n}\geq Cb_{n}$ ($a_{n}\leq Cb_{n}$)., and even after 95% reduction, each iteration would still require at least $10^{8}$ operations and memory storage. For the update step, the computational cost for the cluster center depends on the pre-specified cost function $c(\cdot,\cdot)$. The KL divergence considered in Zhang et al. (zhang2023gaussian, ) suffers from (a) significant overhead of computing $O(\rho s^{2}\log n)$ covariance matrices inversions, and (b) numerical instability due to small eigenvalues of covariance matrices. To overcome this challenge, we introduce two key optimizations designed for 3DGS:
  • Blockwise GMR via KD-Tree: To improve computational and memory efficiency during training, we partition the scene into spatially localized blocks and perform GMR independently within each block. As demonstrated in Remark 1, this blockwise approach yields significant computational savings. While both KD-trees (bentley1975multidimensional, ) and Octrees (meagher1982geometric, ) are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages. First, it produces more balanced partitions across regions. Second, it avoids unnecessary subdivisions in sparse regions that would waste computational resources.
@@
  • Efficient Cost Function: We introduce a novel cost function that overcomes the limitations of the KL divergence used in (zhang2023gaussian, ) by being computationally efficient without sacrificing approximation quality. Our proposed divergence is:
@@
    which offers three significant advantages: First, it preserves distributional similarity, as Gaussian distributions are uniquely determined by their mean and covariance. Second, the assignment step requires only efficient vector and matrix norm computations. Third, the update step simplifies to calculating weighted averages, thereby avoiding the computationally expensive covariance matrix inversions required by the KL divergence:
    $\bar{\mu}_{j}^{(t)}=\frac{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}}{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\frac{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}}{\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}}.$
@@
4:1. Build KD-tree from Gaussians $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$ $\triangleright$ $\mathcal{O}(nd\log n)$
5:2. For each leaf block ${\mathcal{B}}_{k}$: $\triangleright$ $\mathcal{O}(nmT/2^{d}$)
@@
[…]
@@
3.2.2 Appearance Optimization
Following geometric compaction, we initialize the appearance attributes (opacity and color) of the compacted Gaussian primitives. For each primitive in the reduced mixture, we assign the appearance parameters from its closest counterpart in the original Gaussian mixture. Using these initial values, we then optimize the appearance attributes through backpropagation within the standard 3DGS training pipeline used in the first stage. We optimize the opacity instead of directly using the values from the GMR algorithm because its output weights do not necessarily satisfy the constraint that opacity must be between 0 and $1$. Fine-tuning the opacity leads to better visualization performance.